X - 15 = -1 kindly suggest.... how to solve this algebra question ...
Found 2 solutions by Alan3354, jim_thompson5910:
Answer by Alan3354(69443)

X - 15 = -1 kindly suggest.... how to solve this algebra question ...
----------
It's not an "algebra question."
It's not a question at all.
It's an equation.
==========
X - 15 = -1
Solve for x.
----
Add 15 to each side.

Answer by jim_thompson5910(35256)

Add 15 to both sides to get

x-15 = -1

x-15+15 = -1+15

x = 14